As the number of students grew, the school hired more teachers. Write an equation to show the relationship between teachers and students from the chart. S=students
t=teachers
Students 100 200 300 400
Teachers 7 12 17 22
I can see that the number of teachers increases by 5 with every additional 100 students but for the first 100 students there are 7 teachers. I can not figure this out. I know that it has to be simple and I just can not see it. Thanks for the help.

Answer by kellytn78(2)

The equation is:
t = s/100 * 5 + 2
for example: 600/100 * 5 + 2 = t
6 * 5 + 2 = t
30 + 2 = t
32 = t
1000/100 * 5 + 2 = t
10 * 5 + 2 = t
50 + 2 = t
52 = t